OpenStudy (anonymous):

solve using the multiplication principle. -5/8 x = -9/10

OpenStudy (anonymous):

multiply both sides by -1, then multiply both sides by 8 and finally multiply both sides by 1/5

OpenStudy (anonymous):

like this? 5/40 x = 9/50

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

post what you got when you multiplied both sides by -1

OpenStudy (anonymous):

5/8 x = 9/10

OpenStudy (anonymous):

then post what you got after multiplying both sides by 8

OpenStudy (anonymous):

5x = 72/10

OpenStudy (anonymous):

then post what you get when you multiply both sides by 1/5

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

x = 72/50

OpenStudy (anonymous):

there you go. you can cancel out 2 in the numerator and denominator to get x = 36/25

OpenStudy (anonymous):

wow, i did it! thx!

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

what is 2/3 times 3/2?

OpenStudy (anonymous):

6/6 =1?

OpenStudy (anonymous):

what is 4/3 times 3/4?

OpenStudy (anonymous):

12/12 = 1

OpenStudy (anonymous):

so what is -5/8 times -8/5

OpenStudy (anonymous):

40/40 = 1

OpenStudy (anonymous):

so -5/8x * -8/5 = -9/10*-8/5 x = 9*8/5*10 x = 72/50 = 36/25

OpenStudy (anonymous):

the lesson here is if you multiply a number n by 1/n, you will always get 1. 1/n is called the multiplicative identity of n

OpenStudy (anonymous):

oh ok, i see now. great! thx!